OpenStudy (anonymous):

would points 3,4 and 7,7 be parallel to 0,0 and -6,8

OpenStudy (texaschic101):

one way to find out is to find the slopes of your points....parallel lines will have the same slope. slope = (y2 - y1) / (x2 - x1) (3,4)...x1 = 3 and y1 = 4 (7,7) ..x2 = 7 and y2 = 7 now we sub slope = (7 - 4) / (7 - 3) slope = 3/4 now do the same thing with the other 2 points... slope = (y2 - y1) / (x2 - x1) (0,0)...x1 = 0 and y1 = 0 (-6,8)..x2 = -6 and y2 = 8 now we sub slope = (8 - 0) / (-6 - 0) slope = 8/-6 slope = - 4/3 as you can see the slopes are not the same. Therefore, the lines are not parallel. In fact, they are negative reciprocals of each other making them perpendicular.